1.   Calculate the kinetic energy and momentum of a proton traveling 2.90 𝑥 108 𝑚/𝑠.



1
Expert's answer
2022-04-04T09:05:50-0400

The relativistic kinetic energy of a proton is

"K=\frac{1}{\sqrt(1-(\frac{v}{c})^2)} mc^2"

"=\frac{1}{\sqrt(1-(\frac{2.90 \times10^8}{3 \times10^8})^2)} (938 MeV)
 =3663.5 MeV"

The relativistic momentum of a proton is

"p=\frac{1}{\sqrt(1-(\frac{v}{c})^2)} mv"

"=\frac{1}{\sqrt(1-(\frac{2.90 \times10^8}{3 \times10^8})^2)} (938 MeV/c^2)(\frac{2.90 \times10^8}{3 \times10^8})c"

"=35141.5 MeV\/c"